The Perimeter of a triangle is 90 inches. The longest side is 20 inches longer than the shortest side and 10 inches longer than the remaining side.
a)Let x, y, and z be the lengths of the three sides from the longest to smallest. Write a system of three linear equations whose solution gives the lengths of each side.
b) Solve the system
c) Check your solution

Answer by mananth(16949)

x longest
y third side
z = smallest side
perimeter = 90

x=z+20
z=x-20.................(1)
x=y+10
y=x-10.................(2)
x+y+z=92
x+x-10+x-20=90
3x-30=90
3x=120
x= 40 in the longest side
y=x-10=30 in
z=x-20=20in
m.ananth@hotmail.ca